# Quantitative agreement between experiment and theory for Vibrational Circular Dichroism enhanced by electronically excited states

### v1

@@ Abstract
Intensity enhancement in vibrational circular dichroism (VCD) arises in open-shell transition metal complexes from coupling between low-lying excited states (LLESs) and ground-state vibrational modes. In this work we apply Nafie’s vibronic coupling theory to M(II)-(-)-sparteine-Cl2 to investigate these enhancement effects. We show that the VCD intensity is extremely sensitive to the excitation energies that neither time-dependent density functional theory (TDDFT) nor state-averaged complete active space self consistent field (SA-CASSCF) calculations can predict with sufficient accuracy. We argue that instead of using more accurate quantum chemistry methods these excitation energies can be treated as parameters and optimized against experimental spectra. With this approach we obtain simulated VCD similarity scores above 0.4, a threshold considered reliable for absolute configuration assignment.
@@ 1 Introduction
Natural products and other molecules of interest for drug discovery often contain one or more chiral centers that are crucial for their biological activity. This characteristic makes it essential to be able to determine their absolute configuration (AC) or enantiomeric composition 1, 2, 3, 4, 5. In addition to electronic circular dichroism (ECD) spectroscopy, vibrational circular dichroism (VCD) spectroscopy offers a valuable complementary approach for structure determination6. VCD spectra are feature-rich, provide detailed information of the chemical bonds close to the chiral centers, and give insight into intermolecular interactions 7. Many reviews cover both experimental and theoretical VCD methods 3, 8, 2, 4, 9, 10, however, they typically focus on closed-shell systems whose electronic ground state is well separated from excited states. In those cases the VCD spectrum can be efficiently modelled using the magnetic field perturbation (MFP) theory of Stephens11 although the uncertainty in conformer energies can affect the reliability of the stereochemical assignment 12. For molecules with low-lying excited states, as is typically the case in transition metal complexes, the situation is different as the vibrational and electronic degrees of freedom of the molecule are not well-separated. Such chiral molecules are relevant in chiral material science and asymmetric catalysis 13, 14 and there is therefore a growing interest in VCD spectroscopy of molecules containing transition metals 15, 16, 17, 18, 19 that show notably intense VCD peaks.
@@
The phenomenon of enhanced VCD signals in transition metal complexes has been known since 1980, when the first experimental VCD spectra for (--)-sparteine complexes M(II)(sp)Cl2 (M = Zn, Co, Ni) were reported in a narrow region of CH stretching modes 20. Later He et al. extended the measurement to the fingerprint region, where the open-shell Co(II) and Ni(II) have about ten times higher VCD intensity compared to the closed-shell Zn complex 21. Pescitelli and Bari recently published a comprehensive review compiling available data on VCD enhancement and classifying the systems that exhibit this phenomenon 22. They combined the systems with an open-shell metal center into a separate group, which contains mononuclear and multinuclear transition metal complexes, lanthanide complexes, and heme proteins. In most cases, the enhancement can be attributed to the coupling of vibrational and electronic excitations; however, other mechanisms can also come into play  22. Sérgio et al. have reported enhanced VCD spectrum for the ligand attached to paramagnetic Co(II)  23, 24, where these systems show enhanced peaks up to two orders of magnitude higher than the usual VCD peaks. They showed that isostructural Co(III) complexes do not exhibit this intensity amplification, providing experimental evidence that the observed enhancement originates from vibronic coupling with low-lying excited states (LLES). In complexes with Schiff base ligands, a sign reversal effect was observed along with the enhanced VCD signals, and this phenomenon was also attributed to the presence of low-lying excited states in a series of studies by Pescitelli 25, 26, 27. Enhancement due to LLES has also been observed in M(III) complexes: for example, Ru(III)(acac)3 studied by More et al. 28, and a tenfold enhancement was reported for [Co(III)stien(biur)2]NBu4 29.
@@
The enhanced intensity provides two practical benefits. First, the overall sensitivity of VCD improves significantly, for example, Lorenzo et al. applied this mechanism to enable a rapid and sensitive chirality recognition in $\alpha$-amino acids 30. Second, the enhancement is distance-dependent; Domingos et al. demonstrated that amplification occurs locally within specific regions of the molecule, providing a way to “zoom in” on particular parts of a biomolecular system 31. However, in order to be able to use these potential applications, a reliable computational approach is needed to analyze and predict the observed enhanced VCD spectra.
@@
Nafie developed the theoretical description of these phenomena, employing a sum-over-states (SOS) formalism that incorporates coupling between vibrational normal modes and LLES 32. Although Nafie’s VCD vibronic theory  32 provides a rigorous framework for vibronic enhancement, it has not been implemented in major quantum chemical software packages, and to date there have been no published ab initio calculations that successfully reproduce enhanced VCD spectra. Tomeček and Bour suggested an alternative approach and studied Co(II)-(-)(sparteine)Cl2 and bis[(S)-N-(1-phenylethyl)salicylaldiminato]$\Delta$-Co(II) complexes. They employed a Herzberg-Teller expansion of the electronic wave function and incorporated nonadiabatic contributions in a model Hamiltonian, followed by full diagonalization of Hamiltonian on the basis of adiabatic wave functions. Although their approach demonstrated clear amplitude amplification, it was still unable to capture the correct signs for several VCD bands. For example, in Co(sp)Cl2, their calculated spectrum reproduced the overall band shape reasonably well, but the predicted enhancement remained roughly three times smaller than the experimental value. In the case of bis[(S)-N-(1phenylethyl)salicylaldiminato]$\Delta$-cobalt(II), their method could not generate the characteristic monosignate VCD profile observed in the experiment.
@@
In this work, we adopt Nafie’s vibronic coupling theory of VCD and argue that previous attempts to predict enhanced VCD failed because existing quantum chemistry methods cannot describe low-lying excited states with the required accuracy. We propose a computational protocol to address these limitations and demonstrate its effectiveness for Me(II)–(--)-sparteine complexes. The article is organized as follows. In Section 2, we review Stephens’ MFP formalism for VCD. Nafie’s vibronic coupling theory is summarized in Section 3. Computational details are provided in Section 4. In Section 5, we present and discuss our results for Co(II) and Ni(II) sparteine complexes. We therein show that, despite the uncertainties in excitation energy prediction by state-of-the-art quantum chemical methods and neglect of Franck-Condon effects, reliable identification of absolute chirality is possible. This opens the way for application of VCD for structure determination of a much broader range of compounds than has been possible with MFP.
@@ 2 Magnetic Field Perturbation theory (MFP)
In Stephens MFP formalism for VCD 11, the VCD rotational strength $R_{i}$ is written as 11, 33, 34,
@@
where $i$ is the $i^{th}$ normal mode, $\mathbf{E}^{\textup{tot}}$ and ${\mathbf{M}}^{\textup{tot}}$ are the total electric and magnetic dipole transition moments, respectively. The $\mathbf{E}^{\textup{tot}}$ and ${\mathbf{M}}^{\textup{tot}}$ vectors are defined as follows,
$E^{\textup{tot}}_{\beta,i}={\Bigg{(}{\frac{\hbar}{{\omega}_{i}}}\Bigg{)}}^{\frac{1}{2}}\sum_{\lambda\alpha}P_{\alpha\beta}^{\lambda}S_{\lambda\alpha,i},$ (2)
@@
where $\mathbf{S}$ is the transformation matrix from Cartesian ($\lambda\alpha$) to normal ($i$) nuclear coordinates, $\mathbf{P}$ and $\mathbf{A}$ are the atomic polar tensor (APT) and the atomic axial tensor (AAT), respectively. $\omega_{i}$ is the harmonic angular frequency of the $i^{th}$ mode, $\lambda$ is used to enumerate the nuclei and $\alpha$ and $\beta$ indicate Cartesian components of either the nuclear displacement vectors or the electromagnetic field, respectively. The APT and AAT tensors can be written as sum of electronic and nuclear contributions. The APT tensor is then given as,
$P_{\alpha\beta}^{\lambda}=E_{\alpha\beta}^{\lambda}+N_{\alpha\beta}^{\lambda},$ (4)
with the electronic contribution
$E_{\alpha\beta}^{\lambda}={\frac{\partial\langle\Psi_{g}(\mathbf{R})|\mu_{\beta}|\Psi_{g}(\mathbf{R})\rangle}{\partial R_{\lambda\alpha}}\Bigg{|}}_{\mathbf{R}=\mathbf{R}_{0}},$ (5)
and the nuclear contribution
@@
where $\mathbf{I}^{\lambda}$ is the electronic contribution and $\mathbf{J}^{\lambda}$ is the nuclear contribution. The electronic contribution is given as the overlap of the perturbed wave functions with respect to nuclear displacements and with respect to magnetic field,
$I_{\alpha\beta}^{\lambda}={\Bigg{\langle}\frac{\partial\Psi(\mathbf{R},\mathbf{B})}{\partial R_{\lambda\alpha}}\Bigg{|}\frac{\partial\Psi(\mathbf{R},\mathbf{B})}{\partial B_{\beta}}\Bigg{\rangle}\Bigg{|}}_{\mathbf{R}=\mathbf{R}_{0},\mathbf{B}=0},$ (8)
@@
$J_{\alpha\beta}^{\lambda}=i\frac{eZ_{\lambda}}{4\hbar c}\sum_{\gamma}{\epsilon}_{\alpha\beta\gamma}R_{\lambda\gamma}^{0}$ (9)
@@ 3 Nafie’s vibronic theory of VCD with LLES
Stephens’ efficient MFP formalism has been implemented in many of the quantum chemistry programs 11, 33, 5 but does not account for the vibronic coupling with LLESs. Equations to express this coupling in terms of additional LLESs contributions to the APT and the AAT tensors have been derived by Nafie 32 and will be employed here. We refer to this paper for the full derivation and will only briefly discuss the SOS formulation of the APT and the AAT tensors.
@@
On the RHS of Eq. 10, ${\Psi_{g}}$ is the ground state electronic wave function and ${\Psi_{e}}$ refers to electronically excited-state wave functions. The frequency corresponding to a vertical electronic excitation energy is denoted as $\omega_{eg}$ while the normal mode frequency is given as $\omega_{a}$. In this formalism, the APT is dependent on the normal mode frequency, but it can be related to the one defined by Stephens by writing Eq. 10 in terms of two contributions,
@@
The first term in the RHS of Eq. 11, is equivalent to the APT defined in the MFP formalism,
@@
where $\mathbf{m}$ is the magnetic dipole operator and we have multiplied the original definition of Nafie by a factor of $\frac{i}{2}$ to be consistent with the one of Stephens8. Eq. 13 can be partitioned as,
@@
and the second term of Eq. 14 represents the enhancement.
@@
The enhancement corrections can be calculated separately and require evaluation of excitation energies, electric-dipole transition, magnetic dipole transition for the LLES as are readily available in most quantum chemistry codes. The required non-adiabatic coupling vectors (NACVs) $\langle\Psi_{e}|\frac{\partial\Psi_{g}}{\partial R_{A\alpha}}\rangle$ can also be obtained from a number of programs.
Nafie’s formulation employs the vertical approximation, in which the ground and excited state potential energy surfaces are assumed to have the same shape and differ only by a constant energy shift. Within this approximation, one assumes $\phi^{a}_{ev}=\phi^{a}_{gv}$ for all vibrational states ($v$) of elecronic ground ($g$) and excited states ($e$). Consequently, overlap integrals of the form $\braket{\phi^{a}_{g0}}{\phi^{a}_{ev}}$ are zero for all transitions except the fundamental $0\rightarrow 0$ transition. As we show in the results section, however, this vertical approximation is not holding for all low-lying excited states of sparteine complexes and can be one of the reasons of deviations of the calculated and observed spectra.
@@ 4 Computational details
We study the low-lying excited states effects on the VCD spectra of the open-shell complexes Co(II)(sp)Cl2 and Ni(II)(sp)Cl2 . The cobalt (d7 electronic configuration) and nickel (d8 electronic configuration) atoms form high-spin open-shell, four-coordinate complexes with the bidentate chiral (-)-sparteine ligand with the absolute configuration of 6R, 7S, 9S, 11S ( Figure 1). They can be compared with analogous zinc complexes where the electronic structure is that of a closed shell configuration with no unpaired spins, significantly higher excitation energies, and therefore negligible enhancement effects. To gain more insight about the electronic structure of the excited state, we also study model complexes with $C_{2v}$ symmetry. These model complexes are constructed by truncating and symmetrizing the optimized sparteine complexes while preserving bond distances around the transition metal.
@@
Figure 1: Co(sp)Cl2 structure optimized on BP86-D3BJ/def2-TZVP level with CPCM treatment of CHCl3 solvent effects (left) and the $C_{2v}$ Co(NH3)2Cl2 model structure.
@@
According to previous studies  35, the sparteine complexes are inherently rigid as a result of the ligand structure itself. To provide an independent verification of this finding, we use CREST  36 within the Amsterdam Modeling Suite (AMS)  37 to generate initial conformers of Zn(sp)Cl2 employing the GFN1-xTB method  38. Conformers within 5 kcal/mol of the lowest energy structure are subsequently reoptimized at the DFT level using BP86 39, 40-D3BJ 41, 42/def2-TZVP with the conductor-like polarizable continuum model (CPCM)  43 to model CHCl3 solvent effects. All the IR and MFP VCD are performed on the same level of theory for closed-shell Zn(sp)Cl2 and open-shell Co(sp)Cl2 and Ni(sp)Cl2 complexes.
@@
We calculate excited states of open-shell complexes using CASSCF and TDDFT, including dipole transition moments (EDTM/MDTM) and non-adiabatic couplings. TDDFT calculations are performed with (ORCA 6.0) 44. We use def2-TZVP 45 basis set with def2/JK 46 auxiliary sets and extract electric and magnetic transition dipole moments (EDTM/MDTM) and non-adiabatic couplings for further enhancement calculations. CASSCF calculations are performed with both ORCA 6.0 and Openmolcas (v23.06) 47. In ORCA, we perform state-averaged (SA-) and state-specific (SS-) CASSCF on sparteine and model $C_{2v}$ complexes with the def2-TZVP basis set and the def2 / JK auxiliary fitting. OpenMolcas is employed specifically for its implementation of non-adiabatic coupling vectors at the SA-CASSCF level; there, we use the ANO-RCC-VTZP basis on the metal  48, N, and Cl atoms  49, and the minimal ANO-RCC-MB basis on C and H.
@@
We calculate enhanced VCD spectra using our developed Python implementation of Nafie’s vibronic coupling theory described in Section 3. To quantify the agreement of the calculated IR and VCD spectra with the observed ones we use Shen  50 simIR and simVCD metrics. To this end, we digitized the experimental IR and VCD spectra for all sparteine complexes from the thesis of He 51 using a Python script. A frequency scaling factor is obtained by maximizing the simIR between the computed and experimental IR spectra over the 950–1500 cm-1 window, and then the same factor is applied to the calculated VCD spectra.

### v2

@@ Abstract
Intensity enhancement in vibrational circular dichroism (VCD) arises in open-shell transition metal complexes from coupling between ground-state vibrational transitions and magnetic dipole-allowed transitions to low-lying excited states (LLESs). In this work we apply Nafie’s vibronic coupling theory to M(II)-(-)-sparteine-Cl2 (M=Zn, Co, Ni) complexes to investigate these enhancement effects. We show that the VCD intensity is extremely sensitive to the excitation energies that neither time-dependent density functional theory (TDDFT) nor state-averaged complete active space self consistent field (SA-CASSCF) calculations can predict with sufficient accuracy. We argue that instead of using more accurate quantum chemistry methods these excitation energies can be treated as parameters and optimized against experimental spectra. With this approach we obtain simulated VCD similarity scores above 0.4, a threshold considered reliable for absolute configuration assignment. The ability to quantitatively reproduce enhanced experimental spectra with computations opens up new research areas, offering amongst else unique possibilities for the study of chiral structure of systems such as transition metal complexes and metalloproteins that so far remained intractable.
@@ 1 Introduction
Natural products and other molecules of interest for drug discovery often contain one or more chiral centers that are crucial for their biological activity. This characteristic makes it essential to be able to determine their absolute configuration (AC) or enantiomeric composition 1, 2, 3, 4, 5. Electronic circular dichroism (ECD) spectroscopy is a spectroscopic method often employed to obtain insight, but its structural information content is relatively limited. Vibrational circular dichroism (VCD) offers in this respect significantly more information.6. VCD spectra are feature-rich, provide detailed information of the chemical bonds close to the chiral centers, and give insight into intermolecular interactions 7. Many reviews cover both experimental and theoretical VCD methods 3, 8, 2, 4, 9, 10, but they typically focus on closed-shell systems for which the electronic ground state is well separated from excited states. In those cases the VCD spectrum can be efficiently modeled using the magnetic field perturbation (MFP) theory of Stephens11 although the uncertainty in conformer energies can affect the reliability of the stereochemical assignment 12.
@@
For chiral molecules with low-lying excited states (LLES), as is typically the case in transition metal complexes, the situation is different. Such molecules are highly relevant in chiral material science and asymmetric catalysis 13, 14 and there is therefore a growing interest in applying VCD spectroscopy to probe their chiral structure  15, 16, 17, 18, 19. Due to the presence of these low-lying electronic states the vibrational and electronic degrees of freedom of the molecule are not well separated, which has been found to lead to significant enhancements of VCD signals. The phenomenon of enhanced VCD signals in transition metal complexes has been known since 1980, when the first experimental VCD spectra for (--)-sparteine complexes M(II)(sp)Cl2 (M = Zn, Co, Ni) were reported in a narrow region of CH stretching modes 20. Later He et al. extended the measurement to the fingerprint region, where the open-shell Co(II) and Ni(II) complexes have about ten times higher VCD intensity compared to the closed-shell Zn complex 21. Pescitelli and Di Bari recently published a comprehensive review compiling available data on VCD enhancement and classifying the systems that exhibit this phenomenon 22. They combined the systems with an open-shell metal center into a separate group, which contains mononuclear and multinuclear transition metal complexes, lanthanide complexes, and heme proteins. In most cases, the enhancement can be attributed to the coupling of vibrational and electronic excitations; however, other mechanisms can also come into play 22. Domingos et al. reported VCD spectra for ligands attached to paramagnetic Co(II)  23, 24 for which enhancements up to two orders of magnitude were observed. They showed that isostructural Co(III) complexes do not exhibit this intensity amplification, providing experimental evidence that the observed enhancement originates from vibronic coupling with LLES. In complexes with Schiff base ligands, a sign reversal effect was observed along with the enhanced VCD signals, and this phenomenon was also attributed to the presence of LLES in a series of studies by Pescitelli 25, 26, 27. Enhancement due to LLES has also been observed in M(III) complexes: for example, Ru(III)(acac)3 studied by More et al. 28, and a tenfold enhancement was reported for [Co(III)stien(biur)2]NBu4 29.
@@
The enhanced intensity provides two practical benefits. First, the overall sensitivity of VCD improves significantly. Arrico et al., for example. applied this mechanism to enable a rapid and sensitive chirality recognition in $\alpha$-amino acids 30. Second, the enhancement is distance-dependent as was shown in the studies of Domingos et al. who demonstrated that amplification occurs locally within specific regions of the molecule, providing a way to “zoom in” on particular parts of a biomolecular system 31. However, in order to be able to use these potential applications, a reliable computational approach is needed to analyze and predict the observed enhanced VCD spectra. As yet, such an approach has been notoriously lacking.
@@
Nafie developed the theoretical description of VCD for molecules with LLES, employing a sum-over-states (SOS) formalism that incorporates coupling between vibrational normal modes and LLES 32. Although Nafie’s VCD vibronic theory  32 provides a rigorous framework for vibronic enhancement, it has not been implemented in major quantum chemical software packages, and to date no ab initio calculations have been published that successfully reproduce enhanced VCD spectra. Tomeček and Bour suggested an alternative approach and studied Co(II)-(-)(sparteine)Cl2 and bis[(S)-N-(1-phenylethyl)salicylaldiminato]$\Delta$-Co(II) complexes. They employed a Herzberg-Teller expansion of the electronic wave function and incorporated nonadiabatic contributions in a model Hamiltonian, followed by full diagonalization of the Hamiltonian on the basis of adiabatic wave functions. Although their approach demonstrated clear enhancements, it was still unable to capture the correct signs for several VCD bands. In Co(sp)Cl2, for example, the calculated spectrum reproduced the overall band shape reasonably well, but the predicted enhancement remained roughly three times smaller than what was observed experimentally. Similarly, for bis[(S)-N-(1phenylethyl)salicylaldiminato]$\Delta$-cobalt(II), the characteristic monosignate VCD profile observed in the experiment could not be reproduced by the calculations.
@@
In this work, we adopt Nafie’s vibronic coupling theory of VCD and argue that previous attempts to predict enhanced VCD failed because existing quantum chemistry methods cannot describe low-lying excited states with the required accuracy. Moreover, we discuss that the commonly accepted assumption that the potential energy surfaces of the LLES do not differ significantly from that of the electronic ground state may not always hold and can influence significantly the predicted spectrum. We propose a computational protocol to address these limitations and demonstrate its effectiveness for Me(II)–(--)-sparteine complexes. The article is organized as follows. In Section 2, we review Stephens’ MFP formalism for VCD. Nafie’s vibronic coupling theory is summarized in Section 3. Computational details are provided in Section 4. In Section 5, we present and discuss our results for Co(II) and Ni(II) sparteine complexes, and compare them with the spectrum predicted for the closed-shell Zn(II) sparteine complex. We therein show that – despite the uncertainties in excitation energy prediction by state-of-the-art quantum chemical methods and neglect of Franck-Condon effects – a reliable identification of absolute chirality is possible. This opens the way for application of VCD as a sensitive tool for structure determination of a much broader range of compounds than has been possible with MFP.
@@ 2 Magnetic Field Perturbation theory (MFP)
In Stephens MFP formalism for VCD 11, the VCD rotational strength $R_{i}$ is written as 11, 33, 34
@@
where $i$ is the $i^{th}$ normal mode, and $\mathbf{E}^{\textup{tot}}$ and ${\mathbf{M}}^{\textup{tot}}$ are the total electric and magnetic dipole transition moments, respectively. The $\mathbf{E}^{\textup{tot}}$ and ${\mathbf{M}}^{\textup{tot}}$ vectors are defined as
$E^{\textup{tot}}_{\beta,i}={\Bigg{(}{\frac{\hbar}{{\omega}_{i}}}\Bigg{)}}^{\frac{1}{2}}\sum_{\lambda\alpha}P_{\alpha\beta}^{\lambda}S_{\lambda\alpha,i}$ (2)
@@
where $\mathbf{S}$ is the transformation matrix from Cartesian ($\lambda\alpha$) to normal ($i$) nuclear coordinates, and $\mathbf{P}$ and $\mathbf{A}$ are the atomic polar tensor (APT) and the atomic axial tensor (AAT), respectively. $\omega_{i}$ is the harmonic angular frequency of the $i^{th}$ mode, $\lambda$ is used to enumerate the nuclei and $\alpha$ and $\beta$ indicate Cartesian components of either the nuclear displacement vectors or the electromagnetic field, respectively. The APT and AAT tensors can be written as a sum of electronic and nuclear contributions. The APT tensor is then given by
$P_{\alpha\beta}^{\lambda}=E_{\alpha\beta}^{\lambda}+N_{\alpha\beta}^{\lambda}$ (4)
with the electronic contribution given by
$E_{\alpha\beta}^{\lambda}={\frac{\partial\langle\Psi_{g}(\mathbf{R})|\mu_{\beta}|\Psi_{g}(\mathbf{R})\rangle}{\partial R_{\lambda\alpha}}\Bigg{|}}_{\mathbf{R}=\mathbf{R}_{0}}$ (5)
and the nuclear contribution by
@@
where $\mathbf{I}^{\lambda}$ is the electronic contribution and $\mathbf{J}^{\lambda}$ the nuclear contribution. The electronic contribution is given as the overlap of the perturbed wave functions with respect to nuclear displacements and with respect to magnetic field
$I_{\alpha\beta}^{\lambda}={\Bigg{\langle}\frac{\partial\Psi(\mathbf{R},\mathbf{B})}{\partial R_{\lambda\alpha}}\Bigg{|}\frac{\partial\Psi(\mathbf{R},\mathbf{B})}{\partial B_{\beta}}\Bigg{\rangle}\Bigg{|}}_{\mathbf{R}=\mathbf{R}_{0},\mathbf{B}=0}$ (8)
@@
$J_{\alpha\beta}^{\lambda}=i\frac{eZ_{\lambda}}{4\hbar c}\sum_{\gamma}{\epsilon}_{\alpha\beta\gamma}R_{\lambda\gamma}^{0},$ (9)
@@ 3 Nafie’s vibronic theory of VCD with LLES
Stephens’ efficient MFP formalism has been implemented in many quantum chemistry programs 11, 33, 5 but does not account for vibronic coupling with LLESs. Equations to express this coupling in terms of additional LLESs contributions to the APT and the AAT tensors have been derived by Nafie 32 and will be employed here. We refer to this paper for the full derivation and will only briefly discuss the SOS formulation of the APT and the AAT tensors.
@@
On the RHS of Eq. 10, ${\Psi_{g}}$ is the ground state electronic wave function and ${\Psi_{e}}$ refers to electronically excited-state wave functions. The frequency corresponding to the vertical electronic excitation energy of state $e$ is denoted as $\omega_{eg}$ while the normal mode frequency is given as $\omega_{a}$. In this formalism, the APT is dependent on the normal mode frequency, but it can be related to the one defined by Stephens by writing Eq. 10 in terms of two contributions,
@@
The first term on the RHS of Eq. 11 is equivalent to the APT defined in the MFP formalism,
@@
where $\mathbf{m}$ is the magnetic dipole operator and we have multiplied the original definition of Nafie by a factor of $\frac{i}{2}$ to be consistent with the one of Stephens8. Eq. 13 can be partitioned as
@@
while the second term of Eq. 14 represents the enhancement.
@@
The enhancement corrections can be calculated separately and require evaluation of excitation energies, electric and magnetic dipole transition moments for the LLES as are readily available in most quantum chemistry codes. The required non-adiabatic couplings (NACs) $\langle\Psi_{e}|\frac{\partial\Psi_{g}}{\partial R_{A\alpha}}\rangle$ can also be obtained from a number of programs.
Nafie’s formulation employs the vertical approximation in which the ground and excited state potential energy surfaces are assumed to have the same shape and to differ only by a constant energy shift. Within this approximation, one has $\phi^{a}_{ev}=\phi^{a}_{gv}$ for all vibrational states (labeled $v$) of normal mode ($a$) of the electronic ground ($g$) and excited states ($e$). Consequently, overlap integrals of the form $\braket{\phi^{a}_{g0}}{\phi^{a}_{ev}}$ are zero for all transitions except the fundamental $0\rightarrow 0$ transition. As we show in the results section, however, this vertical approximation does not hold for all low-lying excited states of sparteine complexes and is one of the reasons for previously observed differences between calculated and observed spectra.
@@ 4 Computational details
We study the effects of low-lying excited states on the VCD spectra of the open-shell complexes Co(II)(sp)Cl2 and Ni(II)(sp)Cl2. The cobalt (d7 electronic configuration) and nickel (d8 electronic configuration) atoms form high-spin open-shell, four-coordinate complexes with the bidentate chiral (-)-sparteine ligand with the absolute configuration of 6R, 7S, 9S, 11S ( Figure 1). They can be compared with analogous zinc complexes where the electronic structure is that of a closed-shell configuration without unpaired spins with significantly higher excitation energies, and therefore negligible enhancement effects. To gain more insight about the electronic structure of the excited states, we also study model complexes with $C_{2v}$ symmetry. These model complexes are constructed by truncating and symmetrizing the optimized sparteine complexes while preserving bond distances around the transition metal.
@@
Figure 1: Co(sp)Cl2 structure optimized on BP86-D3BJ/def2-TZVP level with CPCM treatment of CHCl3 solvent effects (left) and the $C_{2v}$ Co(NH3)2Cl2 model structure (right).
@@
According to previous studies 35, the sparteine complexes are inherently rigid as a result of the ligand structure itself. To provide an independent verification of this finding, we use CREST 36 within the Amsterdam Modeling Suite (AMS) 37 to generate initial conformers of Zn(sp)Cl2 employing the GFN1-xTB method 38. Conformers within 5 kcal/mol of the lowest energy structure are subsequently reoptimized at the DFT level using BP86 39, 40-D3BJ 41, 42/def2-TZVP with the conductor-like polarizable continuum model (CPCM) 43 to model CHCl3 solvent effects. All the IR and MFP VCD calculations are performed on the same level of theory for closed-shell Zn(sp)Cl2 and open-shell Co(sp)Cl2 and Ni(sp)Cl2 complexes.
@@
We calculate excited states of open-shell complexes using CASSCF and TDDFT, including dipole transition moments (EDTM/MDTM) and non-adiabatic couplings. TDDFT calculations are performed with ORCA 6.0 44. We use a def2-TZVP 45 basis set with def2/JK 46 auxiliary sets and extract electric and magnetic transition dipole moments (EDTM/MDTM) and non-adiabatic couplings for further enhancement calculations. CASSCF calculations are performed with both ORCA 6.0 and Openmolcas v23.06 47. In ORCA, we perform state-averaged (SA-) and state-specific (SS-) CASSCF on sparteine and model $C_{2v}$ complexes with the def2-TZVP basis set and the def2/JK auxiliary fitting. OpenMolcas is employed specifically for its implementation of non-adiabatic coupling vectors at the SA-CASSCF level. There we use the ANO-RCC-VTZP basis on the metal 48, N, and Cl atoms  49, and the minimal ANO-RCC-MB basis on C and H.
@@
We calculate enhanced VCD spectra using our developed Python implementation of Nafie’s vibronic coupling theory described in Section 3. To quantify the agreement of the calculated IR and VCD spectra with the observed ones we use Shen 50 simIR and simVCD metrics. To this end, we digitized the experimental IR and VCD spectra for all sparteine complexes from the thesis of He 51 using a Python script. A frequency scaling factor is obtained by maximizing the simIR between the computed and experimental IR spectra over the 950–1500 cm-1 window. The same factor is then applied to the calculated VCD spectra.